Peter goes on a diet to lose weight. Let the input t represent the week of his diet, and let the output W represent Peter's weight in pounds. At the start of his diet (t = 0), Peter weighed 235 pounds. After 6 weeks on his diet, he weighed 217 pounds. Find the linear relationship between t and W.
Thanks in advance to whoever helps me solve this!

Answer by Boreal(15235)

let t= number of weeks
then (0, 235) and (6, 217) are true.
point slope formula is y-y1=m(x-x1), m slope and (x1, y1) a point
slope is -18/6=-3
equation is t-235=-3W
W=-3t+235 units are pounds